Given that 29a031 x 342 = 100900b02 where a and b denote two missing interger, find the value of a+b
Answer by RAY100(1637)

a=5,,,,b=6,,,,,a+b= 11
.
check,,,295031 * 342 = 100900602
.
.
work
.
.,,,,,,,,,,,,29a031__
,,,,,342[ 100900b02
,,,,,,,,,,,684
,,,,,,,,,,,___
,,,,,,,,,,,3250
,,,,,,,,,,,3078
,,,,,,,,,,,____
,,,,,,,,,,,,1720,,,,,,,1720/342=5.03,,,,therefore a=5,,,,,5*342=1710
,,,,,,,,,,,,1710
,,,,,,,,,,,,____
,,,,,,,,,,,,,,,10b
,,,,,,,,,,,,,,,,,0
,,,,,,,,,,,,,,,___
,,,,,,,,,,,,,,,10b0
,,,,,,,,,,,,,,,1026
,,,,,,,,,,,,,,,_____
,,,,,,,,,,,,,,,(b-1-2) 42
,,,,,,,,,,,,,,,,,,,,342
,,,,,,,,,,,,,,,,,________
,,,,,,,,,,,,,,,,,,,,,0
.
therefore b-1-2 =3,,,,,or b= 6